2. Observe that 49 = 4


×


9 + 4 + 9.


(a) Find all other two-digit numbers which are equal to the product of their digits plus the sum of their


digits.


(b) Are there three-digit numbers which are equal to the product of their digits plus the sum of their digits?


Prove your answer.



1
Expert's answer
2021-10-05T23:14:52-0400

(a) Let our two-digit number be "10a+b". Then



"10a+b=a\times b+a+b""a\times b=9a"

Since "a\not=0," we have "b=9."

Two-digit numbers which are equal to the product of their digits plus the sum of their digits: "19, 29, 39, 49, 59, 69, 79, 89, 99."


(b) Let our three-digit number be "100a+10b+c". Then



"100a+10b+c=a\times b\times c+a+b+c""a\times b\times c=99a+9b"

Since "a\not=0," we have "b\not=0, c\not=0."



"b\times c\leq9\times9=81<99"

Then



"a\times b\times c<99a+9b."

Therefore there is no three-digit number which is equal to the product of its digits plus the sum of its digits.